# Zilber’s Trichotomy in Hausdorff Geometric Structures

### v1

@@ 1. Introduction
The main goal of this paper was to complete the proof of Zilber’s Restricted Trichotomy Conjecture for algebraically closed fields. In characteristic 00 this conjecture was proven by the first author in [Cas23]. In order to adapt key analytic ideas from [Cas23] to positive characteristic, we opted, rather than working with formal schemes (as in [HS24]), to work in algebraically closed valued fields (ACVF). In this setting – in complete models – the theory of analytic functions provides suitable analogues of the characteristic 0 statements we need. Our main result is, thus:
@@
Let us briefly explain the terminologies used above. Recall that a $\mathcal{K}$-relic (see [CH24]) is a structure $\mathcal{M}=(M,...)$ such that the universe $M$, and all $\mathcal{M}$-definable subsets of powers of $M$, are interpretable sets in $\mathcal{K}$. If $M\subset K^{n}$ for some $n$, we call $\mathcal{M}$ a definable $\mathcal{K}$-relic.
@@
Lacking a unifying conjecture, research focused on restricting attention to strongly minimal relics of various theories of a geometric nature. In the late 1980s, Zilber and Rabinovich studied ACF-relics (see [Rab93] and references therein), subsuming works of Martin [Mar88] and Marker-Pillay [MP90]. In 2006 Peterzil suggested an analogous conjecture for o-minimal relics, and in [KR16] Kowalski and Randriambololona conjectured an analogue for ACVF.
The present work contributes to all of these conjectures. It gives a positive solution to the conjecture restricted to definable ACVF-relics, and thus to ACF-relics. In residue characteristic 0, the result also covers general (i.e. interpretable) ACVF-relics. Moreover, in a forthcoming paper, the first author uses the results of the present work to solve Peterzil’s conjecture for high-dimensional relics (which, combined with the main result of [HS24], gives Zilber’s trichotomy restricted to strongly minimal relics of Compact Complex Manifolds – a result that may be viewed as generalizing Moosa’s non-standard Riemann existence theorem, [Moo04]).
As the proof of Theorem 1 emerged, it became clear that the arguments could be written in greater generality. This led us to introduce an axiomatic framework – which we call Hausdorff geometric structures – and which allows for uniform treatments of crucial steps in various proofs of instances of the trichotomy. Our main abstract result (Theorem 8.1) states:
@@
Theorem 2 relates to a recent work of Onshuus, Pinzon and the second author, [HOP24], where the problem of interpreting a field in definable ACVF-relics reduces to showing that the relic $\mathcal{M}$ interprets a group locally isomorphic to a local subgroup of either $(K,+)$ or $(K^{*},\cdot)$. As this can only happen if $M$ is (as a $\mathcal{K}$-definable set) 1-dimensional, the remaining problem splits in two: (1) showing that non-local modularity of $\mathcal{M}$ implies that $\dim(M)=1$, and then (2) showing that strongly minimal, non-locally modular, 1-dimensional relics interpret a group with the desired properties. We expect that (at least in characteristic 0) the axiomatic framework of Theorem 2 may suffice (with minor adjustments) to produce a field directly, without going through a group first. However, as this is not needed for our main results, and in order to not further lengthen the paper, we have not extended our axiomatization to cover this part of the argument.
@@ 1.1. Strategy of proof
The proofs of our main theorems are split into three uneven parts. Sections 3 through 8 introduce the axiomatic framework and culminate with the proof of Theorem 2. In Sections 9 and 10, we verify that ACVF satisfies the assumptions of Theorem 2, and thus (in combination with the main result of [HOP24]) deduce Theorem 1. Section 12 is independent of the rest of the paper. It uses techniques introduced in [HHP22] to show that interpretable non-locally modular strongly minimal ACVF0,0-relics definably embed into a power of the valued field sort or the residue field sort (and thus are covered by Theorem 1 and Corollary Corollary).
@@
The main technical novelty in this section is the notion of enough open maps – a condition inspired by the most critical use of the complex analytic topology in [Cas23], as well as one of the critical uses of o-minimal geometry in the o-minimal variant [EHP21] of the restricted trichotomy. It aims to state, without assuming any differential structure, that the intersection of definable curves should be transverse, unless an obvious obstruction prevents it. It is designed to be a common generalization of this phenomenon in both the o-minimal and analytic settings, and as such it is rather technical; but the main point is that certain finite-to-one definable maps are stipulated to be open.
@@
Section 4 begins the study of strongly minimal relics of Hausdorff geometric structures (with enough open maps). At this level of generality (covering such contexts as o-minimal strongly minimal relics, 1-h-minimal strongly minimal relics and more) we can define and prove a key topological property, that we call weak detection of closures. Detection of closures, as a key tool for the study of strongly minimal relics, goes back to [PS01], and plays a crucial role in several subsequent works on various restrictions of Zilber’s trichotomy ([HK08], [EHP21], and most importantly in [Cas23]). In Section 4 we define this notion (Definition 4.8), and in the subsequent section we prove one of the main results of this paper:
@@
As was noted in [Cas23], at this level of generality, it seems somewhat over-optimistic to hope for the relic to detect all closure points of definable sets (in the sense, e.g., that if $X$ is $A$-definable in the relic and $x$ is in the frontier of $X$, then the Morley rank of $x$ over $A$ is less than that of $X$). The key idea, adapted from [Cas23] to our axiomatic setting, is to identify general enough situations when our over-optimistic hopes are in fact fulfilled. The proof of this result follows rather closely the analogous result from [Cas23], but has a few differences that we point out as we go.
@@ 2.1. Geometric Structures
For more details on the basic properties of dimension in geometric structures, we direct the reader to e.g. [AH24, §2].
@@ 2.2. Strongly Minimal Structures
The main advantage of strongly minimal structures (viewed within the larger class of geometric structures) is the notion of stationarity. In the context of geometric structures, a definable set is stationary if it cannot be partitioned into two generic definable subsets. A key feature of stationary sets is that their almost equality classes can be coded: that is, if $X$ is stationary, there is $c\in\mathcal{M}^{eq}$ such that for all $\sigma\in Aut(\mathcal{M})$, $\sigma(c)=c$ if and only if $\sigma(X)$ is almost equal to $X$. This tuple $c$ is unique up to interdefinability, and is called the canonical base of $X$ and denoted $\operatorname{Cb}(X)$. (See [Mar02, §8.2] for more on canonical bases and [Cas23, §2.3] for some relevant applications).
@@
Following [Cas23], we call a plane curve $C\subset M^{2}$ non-trivial if both projections $C\rightarrow M$ are finite-to-one (equivalently, no stationary component of $C$ is almost equal to a horizontal or vertical line). By uniform finiteness, non-triviality is definable in families. We will largely be able to assume that all plane curves considered are non-trivial – but we will clarify this as we go.
We will frequently use intersections of relic-definable families of curves (as well as higher dimensional objects) to study an ambient topology from within the relic. Since, a priori, these curves are not geometric objects, we have to identify some combinatorial properties of families of curves allowing us to study them geometrically. This is discussed extensively in [Cas23, §2], so we will be brief.
@@
Finally, we recall some special types of families of plane curves that were used extensively in [Cas23]. These are thought of as families that have been presented most conveniently and had various types of irregularities removed.
@@ 3.1. The Definition
Our reason for making this choice (i.e. to not require a definable basis) is given by two of the motivating examples. First, in the complex numbers with the analytic topology (considered in [Cas23]), the only definable open sets are the Zariski open sets, which do not form a basis for the analytic topology. Second, in many éz fields (see below for a discussion of these fields), one has only an ind-definable basis (i.e. there is a basis that is a countable union of definable families of open sets). Thus, Definition 3.1 is designed to accommodate these settings. Note, however, that in all other examples we know, the topology does have a definable basis, and this issue is irrelevant.
@@ 3.4. The Open Mapping Property and Enough Open Maps
    Suppose $f:X\rightarrow Y$ is an $A$-definable projection, $x\in X$ and $y=f(x)\in Y$ are both generic over $A$. Let $Z$ be definable over $B$ so that $y\in Z$ is generic over $B$ and $Z\cap U\subset Y\cap U$ for some neighborhood $U$ of $y$. Then $x\in(f^{-1}(Z\cap Y))^{S}$.
@@
The argument in [Cas23] crucially used that the complex field has the open mapping property, which follows from the usual open mapping theorem. Moreover, we will see later (Corollary 9.60) that ACVF has the open mapping property in all characteristics. Now we show:
@@ 4.1. Notational Conventions
Following [Cas23], we make the following notational conventions to distinguish between $\mathcal{K}$ and $\mathcal{M}$:
@@ 4.2. Stating the Main Result
Let us begin our work with $\mathcal{M}$ by stating our first main goal. As stated above, the idea is to show that $\mathcal{M}$ detects $\tau$ in a precise sense. Following [Cas23], we make the following definitions:
@@ 4.3. Coherence
Before moving toward the proof of Theorem 4.10, we briefly study the notion of coherence, which allows us to move between the dim and rk functions smoothly, and in many instances to forgo computations with dim in favor of simpler computations with rk. The material in this subsection is analogous to [Cas23, §3.3].
We begin by noting the following basic facts, which are all easy and can be proven word for word as in [Cas23]:
@@
(1), (2), and (3) are exactly as in [Cas23] (Lemmas 3.16 and 3.17). For (5), let $b$ be any relation of $\operatorname{tp}_{\mathcal{K}}(a/A)$ which is independent from $B$ over $A$, and apply (4).
@@
Let us now begin the proof of Theorem 4.10. Almost all of the proof is independent of the topology on $\mathcal{K}$, and uses only dimension computations in $\mathcal{K}$ and $\mathcal{M}$. In particular, most of the proof can be copied word-for-word from [Cas23]. The key difference occurs in an important geometric step in the proof, where the assumption of enough open maps (Definition 3.25) replaces some application of complex geometry in the original proof.
Rather than duplicating twenty pages of material, we will largely give an outline of the proof, pointing out which parts can be copied and which ones involve the topological axioms. In the former case, we will point the reader to the identical steps in [Cas23] for details. We will then focus almost exclusively on the latter case.
@@ 5.1. The Main Argument
As in [Cas23], we deduce the inductive step of Theorem 4.10 from a different, more technical statement (Proposition 5.4 below), whose proof contains the main geometric idea of the argument as a whole. This proposition will apply to generic members of almost faithful families of non-trivial hypersurfaces. Let us first recall what this means:
@@
As discussed in Section 2.2, we say that an $\mathcal{M}$-definable family $\{H_{t}:t\in T\}$ of $r$-hypersurfaces is almost faithful if for each $t\in T$, there are only finitely many $t^{\prime}\in T$ with $\operatorname{rk}(H_{t}\cap H_{t^{\prime}})=r$. We will (later on) use the existence of such families. Namely, let $H$ be a stationary non-trivial hypersurface. Then, as also discussed in Section 2.2, there are an $\mathcal{M}(\emptyset)$-definable almost faithful definable family of (without loss of generality non-trivial) hypersurfaces $\{H_{t}:t\in T\}$, and an $\mathcal{M}$-generic $\hat{t}\in T$, such that $H$ is almost equal to $H_{\hat{t}}$. See [Cas23, Lemma 2.25] as well as the discussion following Definition 8.11 loc. cit. for details. Moreover, we may assume $T\subseteq M^{n}$ For some $n$, and if $H$ is defined over a coherent parameter (over $\emptyset$), then $\hat{t}$ will be generic in $T$ (i.e. in the sense of $\mathcal{K}$).
Now, the main technical step of Theorem 4.10 (analogous to [Cas23, Proposition 8.15]) is:
@@
Arguing by contradiction, assume that each of (1), (2), and (3) fails. As in [Cas23], we will proceed by studying the pairwise intersections of $H_{\hat{t}}$ with a certain family of curves in $M^{r+1}$. We begin by defining the various objects we will need for the argument.
@@
A minor point of departure from [Cas23] is the definition of $I$. In [Cas23], $I$ also had a $T$-coordinate, so one took triples $(x,t,u)$ with $x\in H_{t}\cap D_{u}$. Our definition essentially prohibits $t$ from varying. This will make the current version of the argument easier, though the role played by $I$ will be the same.
The following facts are all now easy to see. Note that (5) is Lemma 8.21 of [Cas23].
@@
The following facts about $\hat{u}$ are also straightforward (see [Cas23, Lemma 8.23, Lemma 8.24] for the proofs):
@@
Steps I and II have nothing to do with the topology, and are proven entirely by rank and dimension computations. They can be copied word-for-word from [Cas23]. We will not reproduce the details here.
Step III (or rather an analogous statement) is done in [Cas23] using complex geometry. In that setting, II allows one to reduce III to a problem about smooth complex spaces and holomorphic maps, and apply an open mapping theorem from complex analysis. In the abstract setting, we will prove III using the assumption that $(\mathcal{K},\tau)$ has enough open maps. Indeed, most of the required hypotheses are straightforward to verify, with the most difficult ones being given exactly by II.
Step IV is the key topological argument. Roughly, we show using III that $\hat{u}$ belongs to the closure of an $\mathcal{M}(\hat{t}\hat{p})$-definable set of small rank, and conclude using the inductive hypothesis. This argument is largely unchanged from [Cas23], but we still give the details because it is the most crucial step in the proof of Theorem 4.10.
@@
Exactly the same as Lemma 8.25 of [Cas23]. ∎
@@
Copy Lemmas 8.29, 8.34, 8.39, 8.40, 8.41, and 8.44 from [Cas23]. ∎

### v2

@@ 1. Introduction
The main goal of this paper was to complete the proof of Zilber’s Restricted Trichotomy Conjecture for algebraically closed fields. In characteristic 00 this conjecture was proven by the first author in [Cas24b]. In order to adapt key analytic ideas from [Cas24b] to positive characteristic, we opted, rather than working with formal schemes (as in [HS25]), to work in algebraically closed valued fields (ACVF). In this setting – in complete models – the theory of analytic functions provides suitable analogues of the characteristic 0 statements we need. Our main result is, thus:
@@
Let us briefly explain the terminologies used above. Recall that a $\mathcal{K}$-relic (see [CH24b]) is a structure $\mathcal{M}=(M,...)$ such that the universe $M$, and all $\mathcal{M}$-definable subsets of powers of $M$, are interpretable sets in $\mathcal{K}$. If $M\subset K^{n}$ for some $n$, we call $\mathcal{M}$ a definable $\mathcal{K}$-relic.
@@
Lacking a unifying conjecture, research focused on restricting attention to strongly minimal relics of various theories of a geometric nature. In the late 1980s, Zilber and Rabinovich studied ACF-relics (see [Rab93] and references therein), subsuming works of Martin [Mar88] and Marker-Pillay [MP90]. In 2006 Peterzil suggested an analogous conjecture for o-minimal relics, and in [KR16] Kowalski and Randriambololona conjectured an analogue for ACVF. In [Zil14] Zilber applied Rabinovich’s theorem to give a short model theoretic proof of a conjecture in anabelian geometry [BKT10]. In his paper Zilber called for a new, modern, proof of Rabinovich’s theorem (that was never published), preferably one proving the full Restricted Trichotomy for ACF-relics.
Zilber’s challenge, combined with the above conjectures, rekindled interest in the Trichotomy and ultimately led to the present work. Indeed, our results contribute to all of the conjectures mentioned above. First, and most direct, is the positive solution to the conjecture restricted to definable ACVF-relics, and thus to ACF-relics. In residue characteristic 0, the result also covers general (i.e. interpretable) ACVF-relics.
Secondly, these results have applications outside model theory. Recently the first and second author [CH25] used the Trichotomy for ACF-relics to prove a reconstruction theorem for abelian varieties from a sub-variety, expanding and generalizing [Zil14]. Similar techniques give a partial solution to a conjecture of Booher and Voloch [BV21, Conjecture 2.6] extending Zilber’s main result of [Zil14] to generalized Jacobians. We expect that, augmented by known algebro-geometric techniques, this partial result should suffice for providing a complete proof of that conjecture.
As will be explained below in more detail, Zilber’s intention in calling for a modern proof of Rabinovich’s Theorem was to apply the Zariksi Geometries technology. However, Zariski Geometries – despite being a powerful (and essentially the only) tool in certain settings – turned out to be hard to apply to relics. In practice, recovering a Zariski Geometry in a relic raises serious combinatorial problems that seem insurmountable, particularly for higher-dimensional relics.
In view of the above, the last (and possibly most far reaching) contribution of the current paper consists of setting up the new axiomatic framework of Hausdorff Geometric Structures. The aim of this framework is to streamline the proof of Trichotomy-type results – in particular for relics – giving more flexibility in proving crucial ingredients in such works. This new axiomatic framework has already proved applicable beyond the immediate context for which it was developed. In a recent preprint [Cas24a] the first author builds on the results of the present work to prove the higher dimensional case of Peterzil’s conjecture mentioned above. Combined with the results of the present work, this also provides a complete proof of the Trichotomy for relics of Compact Complex Manifolds. Then in [CH24a], the first and second author use similar techniques to prove analogous results for real closed valued fields (and various expansions thereof).
Our main abstract result in the present paper (Theorem 8.1) states:
@@
Theorem 2 relates to a recent work of Onshuus, Pinzon and the second author, [HOP25], where the problem of interpreting a field in definable ACVF-relics reduces to showing that the relic $\mathcal{M}$ interprets a group locally isomorphic to a local subgroup of either $(K,+)$ or $(K^{*},\cdot)$. As this can only happen if $M$ is (as a $\mathcal{K}$-definable set) 1-dimensional, the remaining problem splits in two: (1) showing that non-local modularity of $\mathcal{M}$ implies that $\dim(M)=1$, and then (2) showing that strongly minimal, non-locally modular, 1-dimensional relics interpret a group with the desired properties. We expect that (at least in characteristic 0) the axiomatic framework of Theorem 2 may suffice (with minor adjustments) to produce a field directly, without going through a group first. However, as this is not needed for our main results, and in order to not further lengthen the paper, we have not extended our axiomatization to cover this part of the argument.
@@ 1.1. Strategy of proof
The proofs of our main theorems are split into three uneven parts. Sections 3 through 8 introduce the axiomatic framework and culminate with the proof of Theorem 2. In Sections 9 and 10, we verify that ACVF satisfies the assumptions of Theorem 2, and thus (in combination with the main result of [HOP25]) deduce Theorem 1. Section 12 is independent of the rest of the paper. It uses techniques introduced in [HHP22] to show that interpretable non-locally modular strongly minimal ACVF0,0-relics definably embed into a power of the valued field sort or the residue field sort (and thus are covered by Theorem 1 and Corollary Corollary).
@@
The main technical novelty in this section is the notion of enough open maps – a condition inspired by the most critical use of the complex analytic topology in [Cas24b], as well as one of the critical uses of o-minimal geometry in the o-minimal variant [EHP21] of the restricted trichotomy. It aims to state, without assuming any differential structure, that the intersection of definable curves should be transverse, unless an obvious obstruction prevents it. It is designed to be a common generalization of this phenomenon in both the o-minimal and analytic settings, and as such it is rather technical; but the main point is that certain finite-to-one definable maps are stipulated to be open.
@@
Section 4 begins the study of strongly minimal relics of Hausdorff geometric structures (with enough open maps). At this level of generality (covering such contexts as o-minimal strongly minimal relics, 1-h-minimal strongly minimal relics and more) we can define and prove a key topological property, that we call weak detection of closures. Detection of closures, as a key tool for the study of strongly minimal relics, goes back to [PS01], and plays a crucial role in several subsequent works on various restrictions of Zilber’s trichotomy ([HK08], [EHP21], and most importantly in [Cas24b]). In Section 4 we define this notion (Definition 4.8), and in the subsequent section we prove one of the main results of this paper:
@@
As was noted in [Cas24b], at this level of generality, it seems somewhat over-optimistic to hope for the relic to detect all closure points of definable sets (in the sense, e.g., that if $X$ is $A$-definable in the relic and $x$ is in the frontier of $X$, then the Morley rank of $x$ over $A$ is less than that of $X$). The key idea, adapted from [Cas24b] to our axiomatic setting, is to identify general enough situations when our over-optimistic hopes are in fact fulfilled. The proof of this result follows rather closely the analogous result from [Cas24b], but has a few differences that we point out as we go.
@@ 2.1. Geometric Structures
For more details on the basic properties of dimension in geometric structures, we direct the reader to e.g. [ALH25, §2].
@@ 2.2. Strongly Minimal Structures
The main advantage of strongly minimal structures (viewed within the larger class of geometric structures) is the notion of stationarity. In the context of geometric structures, a definable set is stationary if it cannot be partitioned into two generic definable subsets. A key feature of stationary sets is that their almost equality classes can be coded: that is, if $X$ is stationary, there is $c\in\mathcal{M}^{eq}$ such that for all $\sigma\in Aut(\mathcal{M})$, $\sigma(c)=c$ if and only if $\sigma(X)$ is almost equal to $X$. This tuple $c$ is unique up to interdefinability, and is called the canonical base of $X$ and denoted $\operatorname{Cb}(X)$. (See [Mar02, §8.2] for more on canonical bases and [Cas24b, §2.3] for some relevant applications).
@@
Following [Cas24b], we call a plane curve $C\subset M^{2}$ non-trivial if both projections $C\rightarrow M$ are finite-to-one (equivalently, no stationary component of $C$ is almost equal to a horizontal or vertical line). By uniform finiteness, non-triviality is definable in families. We will largely be able to assume that all plane curves considered are non-trivial – but we will clarify this as we go.
We will frequently use intersections of relic-definable families of curves (as well as higher dimensional objects) to study an ambient topology from within the relic. Since, a priori, these curves are not geometric objects, we have to identify some combinatorial properties of families of curves allowing us to study them geometrically. This is discussed extensively in [Cas24b, §2], so we will be brief.
@@
Finally, we recall some special types of families of plane curves that were used extensively in [Cas24b]. These are thought of as families that have been presented most conveniently and had various types of irregularities removed.
@@ 3.1. The Definition
Our reason for making this choice (i.e. to not require a definable basis) is given by two of the motivating examples. First, in the complex numbers with the analytic topology (considered in [Cas24b]), the only definable open sets are the Zariski open sets, which do not form a basis for the analytic topology. Second, in many éz fields (see below for a discussion of these fields), one has only an ind-definable basis (i.e. there is a basis that is a countable union of definable families of open sets). Thus, Definition 3.1 is designed to accommodate these settings. Note, however, that in all other examples we know, the topology does have a definable basis, and this issue is irrelevant.
@@ 3.4. The Open Mapping Property and Enough Open Maps
    Suppose $f:X\rightarrow Y$ is an $A$-definable projection, and $x\in X$ and $y=f(x)\in Y$ are both generic over $A$. Let $Z$ be definable over $B$ so that $y\in Z$ is generic over $B$ and $Z\cap U\subset Y\cap U$ for some neighborhood $U$ of $y$. Then $x\in(f^{-1}(Z\cap Y))^{S}$.
@@
The argument in [Cas24b] crucially used that the complex field has the open mapping property, which follows from the usual open mapping theorem. Moreover, we will see later (Corollary 9.60) that ACVF has the open mapping property in all characteristics. Now we show:
@@ 4.1. Notational Conventions
Following [Cas24b], we make the following notational conventions to distinguish between $\mathcal{K}$ and $\mathcal{M}$:
@@ 4.2. Stating the Main Result
Let us begin our work with $\mathcal{M}$ by stating our first main goal. As stated above, the idea is to show that $\mathcal{M}$ detects $\tau$ in a precise sense. Following [Cas24b], we make the following definitions:
@@ 4.3. Coherence
Before moving toward the proof of Theorem 4.10, we briefly study the notion of coherence, which allows us to move between the dim and rk functions smoothly, and in many instances to forgo computations with dim in favor of simpler computations with rk. The material in this subsection is analogous to [Cas24b, §3.3].
We begin by noting the following basic facts, which are all easy and can be proven word for word as in [Cas24b]:
@@
(1), (2), and (3) are exactly as in [Cas24b] (Lemmas 3.16 and 3.17). For (5), let $b$ be any relation of $\operatorname{tp}_{\mathcal{K}}(a/A)$ which is independent from $B$ over $A$, and apply (4).
@@
Let us now begin the proof of Theorem 4.10. Almost all of the proof is independent of the topology on $\mathcal{K}$, and uses only dimension computations in $\mathcal{K}$ and $\mathcal{M}$. In particular, most of the proof can be copied word-for-word from [Cas24b]. The key difference occurs in an important geometric step in the proof, where the assumption of enough open maps (Definition 3.25) replaces some application of complex geometry in the original proof.
Rather than duplicating twenty pages of material, we will largely give an outline of the proof, pointing out which parts can be copied and which ones involve the topological axioms. In the former case, we will point the reader to the identical steps in [Cas24b] for details. We will then focus almost exclusively on the latter case.
@@ 5.1. The Main Argument
As in [Cas24b], we deduce the inductive step of Theorem 4.10 from a different, more technical statement (Proposition 5.4 below), whose proof contains the main geometric idea of the argument as a whole. This proposition will apply to generic members of almost faithful families of non-trivial hypersurfaces. Let us first recall what this means:
@@
As discussed in Section 2.2, we say that an $\mathcal{M}$-definable family $\{H_{t}:t\in T\}$ of $r$-hypersurfaces is almost faithful if for each $t\in T$, there are only finitely many $t^{\prime}\in T$ with $\operatorname{rk}(H_{t}\cap H_{t^{\prime}})=r$. We will (later on) use the existence of such families. Namely, let $H$ be a stationary non-trivial hypersurface. Then, as also discussed in Section 2.2, there are an $\mathcal{M}(\emptyset)$-definable almost faithful definable family of (without loss of generality non-trivial) hypersurfaces $\{H_{t}:t\in T\}$, and an $\mathcal{M}$-generic $\hat{t}\in T$, such that $H$ is almost equal to $H_{\hat{t}}$. See [Cas24b, Lemma 2.25] as well as the discussion following Definition 8.11 loc. cit. for details. Moreover, we may assume $T\subseteq M^{n}$ For some $n$, and if $H$ is defined over a coherent parameter (over $\emptyset$), then $\hat{t}$ will be generic in $T$ (i.e. in the sense of $\mathcal{K}$).
Now, the main technical step of Theorem 4.10 (analogous to [Cas24b, Proposition 8.15]) is:
@@
Arguing by contradiction, assume that each of (1), (2), and (3) fails. As in [Cas24b], we will proceed by studying the pairwise intersections of $H_{\hat{t}}$ with a certain family of curves in $M^{r+1}$. We begin by defining the various objects we will need for the argument.
@@
A minor point of departure from [Cas24b] is the definition of $I$. In [Cas24b], $I$ also had a $T$-coordinate, so one took triples $(x,t,u)$ with $x\in H_{t}\cap D_{u}$. Our definition essentially prohibits $t$ from varying. This will make the current version of the argument easier, though the role played by $I$ will be the same.
The following facts are all now easy to see. Note that (5) is Lemma 8.21 of [Cas24b].
@@
The following facts about $\hat{u}$ are also straightforward (see [Cas24b, Lemma 8.23, Lemma 8.24] for the proofs):
@@
Steps I and II have nothing to do with the topology, and are proven entirely by rank and dimension computations. They can be copied word-for-word from [Cas24b]. We will not reproduce the details here.
Step III (or rather an analogous statement) is done in [Cas24b] using complex geometry. In that setting, II allows one to reduce III to a problem about smooth complex spaces and holomorphic maps, and apply an open mapping theorem from complex analysis. In the abstract setting, we will prove III using the assumption that $(\mathcal{K},\tau)$ has enough open maps. Indeed, most of the required hypotheses are straightforward to verify, with the most difficult ones being given exactly by II.
Step IV is the key topological argument. Roughly, we show using III that $\hat{u}$ belongs to the closure of an $\mathcal{M}(\hat{t}\hat{p})$-definable set of small rank, and conclude using the inductive hypothesis. This argument is largely unchanged from [Cas24b], but we still give the details because it is the most crucial step in the proof of Theorem 4.10.
@@
Exactly the same as Lemma 8.25 of [Cas24b]. ∎
@@
Copy Lemmas 8.29, 8.34, 8.39, 8.40, 8.41, and 8.44 from [Cas24b]. ∎